OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt - 16t^2 where h represents height in feet and r represents the rate in feet per second.

OpenStudy (anonymous):

We are assuming that he golfs the ball off of the ground and the initial height is 0?

OpenStudy (anonymous):

h(t)=-16t^2+96t -16t^2+96t=0 -16t(t-6)=0 t=0,6 6 seconds after he hits it, the ball hits the ground, assuming he golfed it off the ground which is unusual